# Weather Estimation for Integrated Sensing and Communication

### v1

@@ Abstract
One of the key features of sixth generation (6G) mobile communications will be integrated sensing and communication (ISAC). While the main goal of ISAC in standardization efforts is to detect objects, the byproducts of radar operations can be used to enable new services in 6G, such as weather sensing. Even though weather radars are the most prominent technology for weather detection and monitoring, they are expensive and usually neglect areas in close vicinity. To this end, we propose reusing the dense deployment of 6G base stations for weather sensing purposes by detecting and estimating weather conditions. We implement both a classifier and a regressor as a convolutional neural network trained across measurements with varying precipitation rates and wind speeds. We implement our approach in an ISAC proof-of-concept, and conduct a multi-week experiment campaign. Experimental results show that we are able to jointly and accurately classify weather conditions with accuracies of $99.38\%$ and $98.99\%$ for precipitation rate and wind speed, respectively. For estimation, we obtain errors of $1.2$ mm/h and $1.5$ km/h, for precipitation rate and wind speed, respectively. These findings indicate that weather sensing services can be reliably deployed in 6G ISAC networks, broadening their service portfolio and boosting their market value.
@@ I Introduction
One of the most prominent and distinctive features of sixth generation (6G) mobile communication systems will be integrated sensing and communication (ISAC). This technology leverages existing communication infrastructure to perform tasks beyond data transmission, such as drone detection, intruder detection, as well as traffic monitoring [11]. Unlike in positioning technologies, where targets actively send and receive signals, sensing targets remain entirely passive. Beyond the aforementioned target detection-focused use cases, ISAC can also be used to monitor weather conditions in the cellular network coverage.
@@ I-B Related Work
Techniques used for rain classification and estimation include single-layer perceptrons [3], support vector machines [10], convolutional neural networks (CNNs) [10], recurrent neural networks [9], long short-term memorys [10], and gated recurrent units [10]. Additionally, some studies focus on creating rainfall maps [4, 1, 14]. In contrast to these approaches, our method does not rely on microwave static links. Instead, we utilize existing hardware deployed for cellular networks, used for standard communication between next-generation node B-radio units (gNB-RUs) and user equipments (UEs). While we also estimate precipitation rates, our approach differs by leveraging CNNs to process CSI-based periodograms, which are images over range and Doppler shift (speed). These periodograms enables us to perform joint estimation of precipitation rate and wind speed. Furthermore, the aforementioned methods rely on signal attenuation solely from unobstructed fixed links where LoS is assumed, i.e., they are not robust to cluttered environments with multiple obstructions and reflections.
@@ II-B CSI Estimation
To estimate weather conditions, we must first provide an estimate of the CSI matrix. The CSI matrix describes the signal propagation between Tx and Rx and captures reflections from multiple scatterers such as buildings, people, drones, and cars. Depending on the ISAC use case, these scatterers might be considered targets or clutter components. Moreover, the CSI matrix also encodes weather phenomena due to multiple effects such as attenuation and reflections caused by weather conditions. These scatterers and weather phenomena affect the Tx OFDM frame by introducing Doppler shifts and propagation delays within the channel between Tx and Rx. For sensing purposes, the Tx OFDM frame $\mathbf{T}_{s}$ serves only as a reference signal. We estimate the channel by the element-wise division of the Rx OFDM frame $\mathbf{R}_{s}$ by the Tx OFDM frame $\mathbf{T}_{s}$ as
@@ II-C Clutter Removal
The goal of clutter removal is to suppress environmental components irrelevant to the sensing task. For weather detection and estimation, we are solely interested in analyzing the effects of raindrops, dust particles, and snowflakes, while attenuating the contributions from the surrounding environment (e.g. buildings). For clutter removal, we use the clutter removal with acquisitions under phase noise (CRAP) algorithm [7]. We refer to the CSI matrix after CRAP as $\hat{\mathbf{H}}_{s}\in\mathbb{C}^{N_{\text{pad}}\times M_{\text{pad}}}$. This algorithm requires a calibration matrix, which we compute based on measurements of the previous day with the lowest precipitation rate and wind speed.
@@ II-E Feature Engineering
where $n$ and $m$ are the indices of the first and second dimension of the tensor, respectively. After computing the 3D feature tensor $\mathbf{F}_{s}$ from each CSI matrix $\mathbf{H}_{s}$, we normalize the matrices so as to improve convergence on the weather detection and estimation problem. Specifically, we perform a channel-wise normalization: for each channel $\ell=1,\dots,L$, the data is normalized across all samples $S$, $N^{\prime}$ rows, and $M^{\prime}$ columns, such that each channel has zero mean and unit standard deviation. We define the normalized 3D feature tensor as $\bar{\mathbf{F}}_{s}.$ The set of tensors $\left\{\bar{\mathbf{F}}_{s}\right\}_{s\in\mathcal{S}}$ forms the feature dataset.
@@ III-A Task I: Classification
where $\phi_{w,y}$ is the weight of the class $y$ in task $w$, $\left[\hat{\mathbf{x}}_{s,w}\right]_{y}$ is the output of the CNN for sample $s$ in weather metric $w$ and class $y$, and class $y=y_{s,w}$ is the correct label for frame $s$ and weather metric $w$. The predicted label for frame $s$ and weather metric $w$ is then
$\hat{y}_{s,w}=\underset{j=1,\dots,C_{w}}{\max}\!\left(\hat{\mathbf{x}}_{s,w}\right)_{j}.$ (9)
@@ III-B Task II: Regression
A regression problem entails estimating a continuous value based on the input data. Since we have no classes in this case, we assume that $C_{w}=1$, for $w=1,2$. We learn the regression function $g_{\bm{\theta}}$ with a mean squared error (MSE) as
@@
Both loss functions in (8) and (10) rely on reference data for all training samples, i.e. it is a supervised learning problem.
@@ III-C CNN Architecture
  • Max pooling layer with kernel size of $2\times 2$, stride of $2$ and padding of 0.
@@
  • Max pooling layer with kernel size of $2\times 2$, stride of $2$ and padding of 0.
@@
  • Fully connected layer with input size of $16\lfloor N^{\prime}/4\rfloor\lfloor M^{\prime}/4\rfloor$ and output size $C_{w}$, i.e., according to the weather metric $w$.
@@ IV-A PoC Setup
Figs. 4(a), 4(b) show the deployment of the ISAC PoC which consists of:
@@
The Tx and Rx antennas are collocated in a quasi mono-static configuration. The deployment location of the antennas is the roof of a 5 story office building within a city. The antennas are fifth generation (5G) mmWave RUs for Tx and Rx, conforming to the evolved common public radio interface (eCPRI) split 7.2x[12]. The PoC uses 5G $\mu=3$ numerology. Each radio frame consists of $M=1120$ OFDM symbols over $N=1584$ subcarriers spaced by $\Delta f=120$ KHz within a bandwidth of 200 MHz. We use a center frequency of $f_{c}=27.6$ GHz, as available in our over-the-air license for our outdoor deployment. For the PoC, we use a pre-recorded OFDM test frame, which is continuously transmitted over the fronthaul link from the SPU to the gNB-RU. Each subcarrier in the test frame is modulated with a random QPSK constellation point. This test frame is emitted by the gNB-RU towards the environment and the signal reflected by the environment is then received by the sniffer, transmitted over fronthaul to the SPU, which then performs CSI and periodogram computation, as described in Sec. II. Further details on the PoC are available in [18].
@@ V Conclusions and Future Work
Our experiments with measurements from the ISAC PoC have demonstrated that we are capable of achieving accuracies of $99.38\%$ and $98.99\%$ for the classification of precipitation rate and wind speed, respectively. We have also shown that by using our method, we can achieve an average estimation error of $1.2$ mm/h and $1.5$ km/h for precipitation rate and wind speed, respectively. These results show good promises for the deployment of weather sensing in future 6G ISAC networks.

### v2

@@ Abstract
One of the key features of sixth-generation (6G) mobile communications will be integrated sensing and communication (ISAC). While the main goal of ISAC in standardization efforts is to detect objects, the byproducts of radar operations can be used to enable new services in 6G, such as weather sensing. Even though weather radars are the most prominent technology for weather detection and monitoring, they are expensive and usually neglect areas in close vicinity. To this end, we propose reusing the dense deployment of 6G base stations for weather sensing purposes by detecting and estimating weather conditions. We implement both a classifier and a regressor as a convolutional neural network trained across measurements with varying precipitation rates and wind speeds. We implement our approach in an ISAC proof-of-concept and conduct a multi-week experiment campaign. Experimental results show that we are able to jointly and accurately classify weather conditions with accuracies of $99.38\%$ and $98.99\%$ for precipitation rate and wind speed, respectively. For estimation, we obtain errors of $1.2$ mm/h and $1.5$ km/h, for precipitation rate and wind speed, respectively. These findings indicate that weather sensing services can be reliably deployed in 6G ISAC networks, broadening their service portfolio and boosting their market value.
@@ I Introduction
One of the most prominent and distinctive features of sixth-generation (6G) mobile communication systems will be integrated sensing and communication (ISAC). This technology leverages existing communication infrastructure to perform tasks beyond data transmission, such as drone detection, intruder detection, as well as traffic monitoring [11]. Unlike in positioning technologies, where targets actively send and receive signals, sensing targets remain entirely passive. Beyond the aforementioned target detection-focused use cases, ISAC can also be used to monitor weather conditions in the cellular network coverage.
@@ I-B Related Work
Techniques used for rain classification and estimation include single-layer perceptrons [3], support vector machines [10], convolutional neural networks (CNNs) [10], recurrent neural networks [9], long short-term memory networks [10], and gated recurrent units [10]. Additionally, some studies focus on creating rainfall maps [4, 1, 14]. In contrast to these approaches, our method does not rely on microwave static links. Instead, we utilize existing hardware deployed for cellular networks, used for standard communication between next-generation node B-radio units (gNB-RUs) and user equipments (UEs). While we also estimate precipitation rates, our approach differs by leveraging CNNs to process CSI-based periodograms, which are images over range and Doppler shift (relative speed). These periodograms enable us to perform joint estimation of precipitation rate and wind speed. Furthermore, the aforementioned methods rely on signal attenuation solely from unobstructed fixed links where LoS is assumed, i.e., they are not robust to cluttered environments with multiple obstructions and reflections.
@@ II-B CSI Estimation
To estimate weather conditions, we must first provide an estimate of the CSI matrix. The CSI matrix describes the signal propagation between Tx and Rx and captures reflections from multiple scatterers such as buildings, people, drones, and cars. Depending on the ISAC use case, these scatterers might be considered targets or clutter components. Moreover, the CSI matrix also encodes weather phenomena due to multiple effects, such as attenuation and reflections caused by weather conditions. These scatterers and weather phenomena affect the Tx OFDM frame by introducing Doppler shifts and propagation delays within the channel between Tx and Rx. For sensing purposes, the Tx OFDM frame $\mathbf{T}_{s}$ serves only as a reference signal. We estimate the channel by the element-wise division of the Rx OFDM frame $\mathbf{R}_{s}$ by the Tx OFDM frame $\mathbf{T}_{s}$ as
@@ II-C Clutter Removal
The goal of clutter removal is to suppress environmental components irrelevant to the sensing task. For weather detection and estimation, we are solely interested in analyzing the effects of raindrops, dust particles, and snowflakes, while attenuating the contributions from the surrounding environment (e.g., buildings). For clutter removal, we use the clutter removal with acquisitions under phase noise (CRAP) algorithm [7]. We refer to the CSI matrix after CRAP as $\hat{\mathbf{H}}_{s}\in\mathbb{C}^{N_{\text{pad}}\times M_{\text{pad}}}$. This algorithm requires a calibration matrix, which we compute based on measurements of the previous day with the lowest precipitation rate and wind speed.
@@ II-E Feature Engineering
where $n$ and $m$ are the indices of the first and second dimensions of the tensor, respectively. After computing the 3D feature tensor $\mathbf{F}_{s}$ from each CSI matrix $\mathbf{H}_{s}$, we normalize the matrices so as to improve convergence on the weather detection and estimation problem. Specifically, we perform a channel-wise normalization: for each channel $\ell=1,\dots,L$, the data is normalized across all samples $S$, $N^{\prime}$ rows, and $M^{\prime}$ columns, such that each channel has zero mean and unit standard deviation. We define the normalized 3D feature tensor as $\bar{\mathbf{F}}_{s}.$ The set of tensors $\left\{\bar{\mathbf{F}}_{s}\right\}_{s\in\mathcal{S}}$ forms the feature dataset.
@@ III-A Task I: Classification
where $\phi_{w,y}$ is the weight of the class $y$ in weather metric $w$, $\left[\hat{\mathbf{x}}_{s,w}\right]_{y}$ is the output of the CNN for sample $s$ in weather metric $w$ and class $y$, and class $y=y_{s,w}$ is the correct label for frame $s$ and weather metric $w$. The predicted label for frame $s$ and weather metric $w$ is then
$\hat{y}_{s,w}=\underset{j=1,\dots,C_{w}}{\max}\!\left[\hat{\mathbf{x}}_{s,w}\right]_{j}.$ (9)
@@ III-B Task II: Regression
A regression problem entails estimating a continuous value based on the input data. Since we have no classes in this case, we assume that $C_{w}=1$, for $w\in\!\left\{1,2\right\}$. We learn the regression function $g_{\bm{\theta}}$ with a mean squared error (MSE) loss as
@@
Both loss functions in (8) and (10) rely on reference data for all training samples, i.e., it is a supervised learning problem.
@@ III-C CNN Architecture
  • Max pooling layer with kernel size of $2\times 2$, stride of $2$, and padding of 0.
@@
  • Max pooling layer with kernel size of $2\times 2$, stride of $2$, and padding of 0.
@@
  • Two independent fully connected layers with input size of $16\lfloor N^{\prime}/4\rfloor\lfloor M^{\prime}/4\rfloor$ and output size $C_{w}$, for $w\in\!\left\{1,2\right\}$.
@@ IV-A PoC Setup
Figs. 4(a), 4(b) show the deployment of the ISAC PoC, which consists of:
@@
The Tx and Rx antennas are collocated in a quasi mono-static configuration. The deployment location of the antennas is the roof of a 5 story office building within a city. The antennas are fifth-generation (5G) mmWave RUs for Tx and Rx, conforming to the evolved common public radio interface (eCPRI) split 7.2x[12]. The PoC uses 5G $\mu=3$ numerology. Each radio frame consists of $M=1120$ OFDM symbols over $N=1584$ subcarriers spaced by $\Delta f=120$ KHz within a bandwidth of 200 MHz. We use a center frequency of $f_{c}=27.6$ GHz, as available in our over-the-air license for our outdoor deployment. For the PoC, we use a pre-recorded OFDM test frame, which is continuously transmitted over the fronthaul link from the SPU to the gNB-RU. Each subcarrier in the test frame is modulated with a random QPSK constellation point. This test frame is emitted by the gNB-RU towards the environment and the signal reflected by the environment is then received by the sniffer, transmitted over fronthaul to the SPU, which then performs CSI and periodogram computation, as described in Sec. II. Further details on the PoC are available in [18].
@@ V Conclusions and Future Work
Our experiments with measurements from the ISAC PoC have demonstrated that we are capable of achieving accuracies of $99.38\%$ and $98.99\%$ for the classification of precipitation rate and wind speed, respectively. We have also shown that by using our method, we can achieve an average estimation error of $1.2$ mm/h and $1.5$ km/h for precipitation rate and wind speed, respectively. These results show good promise for the deployment of weather sensing in future 6G ISAC networks.